OpenStudy (anonymous):

Help me factor this:

OpenStudy (anonymous):

\[x-y+2\sqrt{x}-2\sqrt{y}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[x-y=(\sqrt{x}-\sqrt{y})(\sqrt{x}+\sqrt{y})\]

OpenStudy (anonymous):

\[2\sqrt{x}-2\sqrt{y}=2(\sqrt{x}-\sqrt{y})\]

OpenStudy (anonymous):

Can you solve it now?

OpenStudy (anonymous):

what is the complete factored form

OpenStudy (anonymous):

I started it, you end it

OpenStudy (anonymous):

okay thanks bud

OpenStudy (anonymous):

What answer do you get?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

cause i plugged in 0

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

cause thats what you do when you factor, you plug in 0

OpenStudy (anonymous):

\[(\sqrt{x}-\sqrt{y})(\sqrt{x}+\sqrt{y})+2(\sqrt{x}-\sqrt{y})\]

OpenStudy (anonymous):

exactly and the once you plug in 0 you get the complete factored form

OpenStudy (anonymous):

Take \[(\sqrt{x}-\sqrt{y})\] as common and finish it.

OpenStudy (anonymous):

the complete factored form is 2

OpenStudy (anonymous):

No, there are two factors x and y, they can have any value, so you can't get any one definite answer, there are infinite answer's for this question, which depends on the values of x ans y

OpenStudy (anonymous):

no actually the factored for is 2, cause when you foil out 2, you get back the original form

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Factorization

OpenStudy (anonymous):

exactly thats what i did and the factored form is 2

OpenStudy (anonymous):

Did you go through the link?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In mathematics, factorization (also factorisation in British English) or factoring is the decomposition of an object (for example, a number, a polynomial, or a matrix) into a product of other objects, or factors, which when multiplied together give the original

OpenStudy (anonymous):

Which standard or grade are you in?

OpenStudy (anonymous):

exaclty so 2 is the decomposition of the object, and when it is mulitplied with 56 gives back the original

OpenStudy (anonymous):

What is the original form?

OpenStudy (anonymous):

x-y+2sqrt(x)-2sqrt(y)

OpenStudy (anonymous):

If you multiply 2 with 56 you get 112

OpenStudy (anonymous):

no you get x-y+2sqrt(x)-2sqrt(y)

OpenStudy (anonymous):

56*2=112

OpenStudy (anonymous):

I am not understanding what you're trying to convey, sorry.

OpenStudy (anonymous):

okay